Circular No. 6377

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1996V IN NGC 3644 P. Garnavich, P. Challis, and R. Kirshner, Harvard-Smithsonian Cen-

ter for Astrophysics report that a spectrogram obtained by P. Berlind with the 1.5-m Tillinghast Telescope on Apr. 11.2 UT shows that SN 1996V is a type-Ia supernova, ∼ 2 weeks past maximum. Narrow emission lines from the host galaxy give a recession velocity of 7050 km/s.

COMET C/1995 O1 (HALE-BOPP) D. Jewitt and M. Senay, University of Hawaii; and H. Matthews, Joint

Astronomy Centre, Hilo, and Herzberg Institute of Astrophysics, Ottawa, write: “We observed C/1995 O1 with the 15-m James Clerk Maxwell Telescope on Apr. 8 UT. The 354-GHz J=4-3 transition of hydrogen cyanide was detected with peak antenna temperature 0.09 K, area 0.11 ± 0.01 K km/s, and central velocity −46.4 ± 0.1 km/s. At adopted coma temperatures of 10 and 15 K, the HCN-production rate is 5 and 2 × 10^{26} s^{−1}, respectively. This is to be compared with the production rate Q(CN) = 1.1 × 10^{26} s^{−1} estimated from optical data (IAUC 6361) on Mar. 19. Within the uncertainties due to modelling the production rates and to the time difference between the measurements, we conclude that HCN is a plausible parent of CN, even at the current heliocentric distance of 4.8 AU. The spectrum may

be viewed at http://www.ifa.hawaii.edu/~jewitt/submm.html.”

SUPERNOVA 1996P IN NGC 5335 P. Sicoli and V. Giuliani, Sormano, Italy, report the following precise

position for SN 1996P, obtained with the 0.5-m reflector (+ CCD) on Apr. 7.926 UT: α = 13^{h}52^{m}52^{s}.61, δ = +2^{o}48′30′′.2 (equinox 2000.0).

COMET C/1996 B2 (HYAKUTAKE) Naked-eye m_{1} estimates: Apr. 4.83, 2.6 (A. Diepvens, Balen, Belgium);

5.84, 2.4 (J.-C. Merlin, Le Creusot, France); 6.81, 2.0 (G. Bakos, Balaton, Hungary; 70^{o} tail in p.a. 45^{o}); 7.81, 2.4 (M. Plsek, Lelekovice, Czech Republic); 8.81, 1.9 (R. Haver, Tolfa, Italy); 10.15, 2.2 (R. Keen, Mt. Thorodin, CO; tail 72^{o} long in p.a. 43^{o}); 11.25, 2.8 (S. J. O’Meara, Volcano, HI; tail 65^{o} long in p.a. 42^{o}).

1996 April 11  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green